Suppose the purchase price of a hybrid delivery truck is $90,000, its residual value in four years is certain to be $15,000, and there is no risk that the lessee will default on the lease. Assume that capital markets are perfect and the risk-free interest rate is 6% APR with monthly compounding.

Calculate the monthly lease payments for a four-year lease of the hybrid delivery truck.

(Please show the calculator commands, "FV, PV, I/Y etc.")
Answer #1

Using financial calculator
PV=-90000
FV=15000
I/Y=6%/12=0.5%
N=12*4=48
CPT PMT=$1,836.38